OpenStudy (anonymous):

What is the value of x? 6x -2(x +4) = 4

OpenStudy (camerondoherty):

Hm, first you should add 2 to both sides What do you get?

OpenStudy (anonymous):

Don't you have to do the distributive property with -2 first though..? 6x -2x +2 =4

OpenStudy (anonymous):

6x-2(x+4)=4 1. Distribute 6x-2x-8=4 2. Add common factors 4x-8=4 3. Subtract 8 on both sides 4x=-4 4. Divide by 4 on both sides x=-1

OpenStudy (camerondoherty):

O gosh, I completely read tht wrong, ugh I apologize xc

OpenStudy (anonymous):

x actually equals 3

OpenStudy (anonymous):

plz fan and give me best response

OpenStudy (anonymous):

Oops i distributed wrong :/ sorrrryy

OpenStudy (anonymous):

But those were the right directions (; so i hoped that helped

OpenStudy (camerondoherty):

He did something wrong though Here, let me show you: 6x -2(x +4) = 4 Step 1: You distribute from the parenthases 6x - 2x - 8 = 4 Step 2: Add 8 to both sides 6x - 2x = 12 Step 3: Combine Common factors 4x = 12 Step 4: Divide both sides by 4 And x = 3 c: Do you get it? I hope I made it clear cx

OpenStudy (anonymous):

yeah that is what I said to

OpenStudy (camerondoherty):

Yea but giving away direct answers is against the rules and CoC... cx

OpenStudy (anonymous):

Oh, okay yeah lol I totally understand now! @camerondoherty

OpenStudy (anonymous):

so what

OpenStudy (camerondoherty):

c: